The width of a rectangle is 2 cm less than a side of a square, and the length is 1 cm more than a side of a square. If the area of the rectangle is 28 cm^2, find the length of a side of the square.
Answer by TimothyLamb(4379)

---
x = side of square
y = length of rect
z = width of rect
---
z = x - 2
y = x + 1
yz = 28
---
z = 28/y
---
z = x - 2
z = 28/y
x - 2 = 28/y
y(x - 2) = 28
---
y(x - 2) = 28
y = x + 1
---
(x + 1)(x - 2) = 28
xx - x - 2 = 28
---
xx - x - 30 = 0
---
the above quadratic equation is in standard form, with a=1, b=-1, and c=-30
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -1 -30
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
x = 6
x = -5
---
the negative root does not fit the problem statement, so use the positive root
---
answer:
x = side of square = 6 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php